# Learning to Recover: Dynamic Reward Shaping with Wheel-Leg Coordination for Fallen Robots

### v1

@@ Abstract
Adaptive recovery from fall incidents are essential skills for the practical deployment of wheeled-legged robots, which uniquely combine the agility of legs with the speed of wheels for rapid recovery. However, traditional methods relying on preplanned recovery motions, simplified dynamics or sparse rewards often fail to produce robust recovery policies. This paper presents a learning-based framework integrating Episode-based Dynamic Reward Shaping and curriculum learning, which dynamically balances exploration of diverse recovery maneuvers with precise posture refinement. An asymmetric actor-critic architecture accelerates training by leveraging privileged information in simulation, while noise-injected observations enhance robustness against uncertainties. We further demonstrate that synergistic wheel-leg coordination reduces joint torque consumption by $15.8\%$ and $26.2\%$ and improves stabilization through energy transfer mechanisms. Extensive evaluations on two distinct quadruped platforms achieve recovery success rates up to $99.1\%$ and $97.8\%$ without platform-specific tuning. The supplementary material is available at https://boyuandeng.github.io/L2R-WheelLegCoordination/
@@ I INTRODUCTION
Wheeled-legged robots, with their integrated wheel–leg structure, have demonstrated unique mobility advantages in complex environments, establishing them as key platforms for long-duration tasks such as inspection and exploration[1, 2, 3, 15]. However, unexpected falls resulting from dynamic disturbances (e.g., collisions, slippage) often interrupt missions, and existing systems typically rely on manual intervention for recovery or preplanned recovery motions, which while they can be effective on flat terrains, they are not robust against diverse terrain geometries, severely limiting operational efficiency and autonomy. Thus, breakthroughs in robust and autonomous post-fall recovery are critical to achieving fully autonomous missions [5]
@@
In essence, recovery after fall is a highly complex motion planning and control problem, necessitating precise posture adjustments through multiple discontinuous contact events[9] while the ground specific geometry can impose additional contact uncertainty challenges, which can compromise the successfully execution of the fall recovery actions. For the controller, deriving an optimal or near-optimal sequence of actions is highly challenging. Owing to the highly nonlinear dynamics of wheeled-legged robots, optimization-based methods often depend on simplified system models and precise state estimation. Consequently, most existing work has focused on pure legged recovery[4] mainly demonstrated on flat terrains, leaving the full potential of wheel–leg collaborative recovery largely unexplored. In contrast, recent advances in reinforcement learning have demonstrated remarkable progress in addressing this issue, offering a new technical pathway toward more flexible and robust recovery strategies.
Figure 1: KYON Legged-wheeled Robot model: Wheel and joint coordination for joint reset and Center-of-Mass adjustment during the recovery process. The presented model corresponds to a new robot under development in the Humanoid and Human Centered Mechatronics laboratory at Istituto Italiano di Tecnologia.
@@
In this paper, we explore a widely used learning-based approach to facilitate rapid post-fall recovery for wheeled-legged robots by leveraging their wheels. We employ an asymmetric Proximal policy optimization(PPO) framework, which integrates dynamic exploration of episodes and curriculum learning to train a network. This design mitigates the sensitivity of recovery and smoothness weight parameters while also avoiding an overly conservative policy tendency. The main contributions of this work are summarized as follows:
@@
  • We propose an Episode-based Dynamic Reward Shaping mechanism combined with curriculum learning to balance exploration and policy convergence. This framework enables expansive exploration of diverse recovery maneuvers in early training while progressively refining posture control toward the target configuration, effectively overcoming the limitations of sparse reward designs.
@@
  • We systematically validate the critical role of wheel actuation in reducing joint torque effort and improving stabilization during recovery. This synergy between wheel dynamics and leg adjustments is rigorously quantified through cross-platform experiments, offering new insights into how wheels can actively assist in upright recovery processes.
@@
  • Extensive experimental validation on two distinct wheeled-legged platforms (KYON and Unitree Go2-W[11]) demonstrates high adaptability of the proposed approach to varying hardware configurations, achieving higher success rates and reduced torque usage without platform-specific tuning, underscoring its robustness and scalability.
@@ II RELATED WORK
In reinforcement learning–based methods, relying solely on the final target posture as the primary task reward often leads to sparse feedback, resulting in insufficient exploration and lower success rates. To address this challenge, our study introduces an Episode-based Dynamic Reward Shaping mechanism that balances exploration and convergence. This approach expands the exploration space in the early stages while ensuring that the policy effectively converges to the desired recovery behavior.
Figure 2: (a)-(c) correspond to the initial states of different episodes.
@@ III-A State initialization and rollout
During the execution of various tasks, wheeled-legged robots may fall or even completely roll over; thus, the recovery controller must accommodate diverse initial conditions. To simulate different fallen states, we randomly initialize the robot’s base orientation and joint angles, set the joint torques to zero, and let the robot free-fall from a height of 1.1 m for 2 seconds. This duration is chosen to ensure that the robot fully collapses onto the ground, as shown in the Fig.2, which helps avoid the situation where the dynamics become unrealistic[6]. The diverse set of observed states improves policy generalization and adaptability. Each episode has a fixed duration of 4 seconds and is not terminated early.
@@ III-B Privileged information based Actor-Critic
Actor Observation: The policy network observes the robot’s state, including current and historical readings (with a time interval of 0.01 s) of the base’s linear velocity, angular velocity, orientation, joint positions, joint velocities, and wheel speeds. This temporal information enables the robot to better evaluate its contact state[6]. The actor also receives the previous action taken. Except for the most recent action and joint positions, Gaussian noise is added to the observations to account for state-estimation inaccuracies that arise after a fall. Similar to[5], we apply higher noise levels specifically to joint velocities and the base’s linear and angular velocities.
@@ III-D Dynamic Exploration strategy
1) Dynamic Reward Shaping:In previous research, the post-fall recovery process has often been simplified to a “final pose pursuit” task, typically employing relatively fixed reward or penalty terms based on joint angles, body height, or orientation error. However, in transitioning from a prone position to standing, a wheeled-legged robot may need to undergo large—and sometimes counterintuitive—rolling or swinging movements. Penalizing only the terminal pose too strongly can restrict exploration of diverse state-action pairs, as illustrated in Fig. 2. To balance free exploration in the early stage with precise posture control in the later stage, we propose Episode-based Dynamic Reward Shaping(DS), formally defined as:
$DS=\left(\frac{a\cdot t}{T}\right)^{k}$ (1)
where $t$ is the current step in an episode, $t\in[0,T]$. $T$ is the total number of steps per episode, and $k$ serves as the growth rate, whereas $a$ is the baseline coefficient. For our experiments, we set $a=\frac{T}{2}$, according to the expected recovery time, $k=3$. The immediate reward at each time step is given by:
$r_{t}=DS\cdot R\left(s_{t},a_{t}\right)$ (2)
Here, $R(s_{t},a_{t})$ is the original baseline reward function. When the $t$ is close to 0, the agent can attempt substantial posture adjustments, and any successful state-action pairs are retained and reinforced in the policy network. As $t$ approaches the maximum length of the episode, the emphasis is progressively increased on the refinement of the final target pose. This incremental mechanism promotes action diversity in the early stage of each episode and ensures the stability of the standing posture in the later stage.
@@
where $\beta$ is the learning rate. We multiply the auxiliary reward by this weight and update it before each environment reset. This approach ensures rapid acquisition of task-related behaviors in the early stages and smoother motion in the later stages, striking a balance between efficiency and feasibility.
@@
body collision $\sum_{b\in B}\|\lambda_{b}[t]\|^{2}$ $-5\times 10^{-2}$
action rate $\sum_{t}(\mathbf{a}[t]-\mathbf{a}[t-1])^{2}$ $-1\times 10^{-2}$
joint velocity $\sum_{j}\dot{\mathbf{q}_{j}}^{2}$ $-2\times 10^{-2}$
torques $\sum_{j}\tau_{j}^{2}$ $-2.5\times 10^{-5}$
acceleration $\sum_{j}\ddot{\mathbf{q}_{j}}^{2}$ $-2.5\times 10^{-7}$
wheel velocity $\sum_{k}\dot{\mathbf{q}_{k}}^{2}$ $-2\times 10^{-2}$
Figure 4: Recovery processes under the same initial posture using two different strategies. (a)–(d) shows the DS strategy successfully recovery. (e)–(h) illustrates the baseline strategy, which adjusts joint positions and base height but fails to optimize base orientation, becoming trapped in a local optimum and not fully recovering.
@@ III-E Reward function
$r_{t}=$ $\,DS\cdot r_{joint\_pos}+r_{base\_height}+r_{base\_ori}$
$+CW\cdot\big{(}r_{collision}+DS\cdot(r_{action\_rate}$
$+r_{joint\_vel}+r_{torque}+r_{joint\_acc}))$
@@ IV-A Experimental Setup
All experiments are conducted in the Isaac Sim simulator using KYON and Unitree Go2-W[11] robot models. To enhance policy robustness, extensive domain randomization is applied: the robot’s base mass is perturbed by ±$\pm 5.0kg$ at the start of each episode, link masses vary by ±$\pm 10\%$ from their nominal values. Moreover, contact dynamics are randomized by uniformly sampling static friction coefficients from $[0.7,1.3]$ for all robot bodies, while actuator stiffness and damping are perturbed by ±$\pm 15.0$ and ±$\pm 2.0$ respectively. Additionally, action noise characterized by a zero mean and a standard deviation of $0.02$ is incorporated.
@@
Each episode lasts 4 seconds (at a control rate of 100Hz) and terminates solely upon timeout. The training framework employs an asymmetric PPO algorithm with a learning rate of $0.001$ and a discount factor of $\gamma=0.99$. Furthermore, curriculum learning is integrated, beginning with a difficulty factor of $0.3$ that decays exponentially at a rate of $0.968$ to progressively enhance policy exploration.
IV-B Baseline Comparison on Recovery Success Rate
@@
To verify the effectiveness of the Episode-based Dynamic Reward Shaping (DS), we compared the recovery performance of the DS-Policy with that of a baseline under an identical reward function and hyperparameter framework. Recovery was defined as achieving, at the final time step of an episode, a base height exceeding $0.42m$, a deviation of the joint angle from the default configuration of no more than $0.5rad$, and a maximum joint velocity below $0.1$ rad/s. Testing on 2048 randomly generated initial conditions, the DS-Policy achieved a success rate of $99.1\%$, representing a $2.7$ percentage point improvement over baseline $96.4\%$.
@@
Further analysis revealed that the failure cases for the baseline were predominantly associated with scenarios in which the front leg joint angles were minimally randomized, shown in Fig4. To illustrate this, we selected a representative initial state from one of the baseline’s failure cases and used it as the starting state for the DS-Policy to compare the resulting action sequences. The experiments indicate that, due to the penalty imposed on joint angle deviations, the baseline suppressed front leg motions and attempted to adjust the base height solely through the hind legs, ultimately becoming trapped in a local optimum.
@@
This phenomenon is directly related to differences in the exploration capacity of the action space. Principal Component Analysis (PCA) was applied to the action sequences from the 2048 experiments, and the resulting visualization Fig.5 shows that the variance along the principal components for the DS-Policy ($\sigma^{2}_{PC1}=20.91$, $\sigma^{2}_{PC2}=9.72$) is significantly higher than that for the baseline ($\sigma^{2}_{PC1}=17.33$, $\sigma^{2}_{PC2}=5.22$). These findings confirm that the dynamic reward shaping mechanism, by expanding the exploration boundaries and generating less conservative action sequences, enhances recovery robustness in complex scenarios.
@@
Figure 5: PCA-based comparison of single-episode action distributions for DS-policy and baseline-policy across 2048 environments.
@@ IV-C Wheel-Assisted Recovery
To evaluate the dynamic contribution of wheeled actuation during recovery, we compared the performance differences between a wheel–leg collaborative mode and a pure legged mode. In the experimental setup, the pure legged group was configured by modifying the robot description file to change the wheel joints from continuously actuated to fixed constraints, and by removing the wheel speed penalty term to eliminate policy bias; all other hyperparameters and reward functions were kept identical to those of the wheel–leg collaborative group. Testing across 2048 random initial states demonstrated that the wheel–leg collaborative mode significantly enhanced recovery stability and optimized joint torque distribution. As shown in Fig.6, although both strategies’ base height trajectories converged to the target height ($0.42m$) within $0.7s$, the wheel–leg collaborative group exhibited lower variance during both the recovery phase ($t<0.7s$) and the stabilization phase ($0.7s<t<1.5s$). In contrast, the pure legged group showed larger fluctuations during the stabilization phase, with a variance of $0.102$ at $t=1s$.

### v2

@@ Abstract
Adaptive fall recovery is a critical capability for the practical deployment of wheeled-legged robots, which uniquely combine the agility of legs with the speed of wheels for rapid and efficient recovery. However, conventional approaches that depend on preplanned motions, simplified dynamics, or sparse rewards often fail to yield robust recovery behaviors. This paper introduces a learning-based framework that unites episode-based dynamic reward shaping with curriculum learning to balance the exploration of diverse recovery maneuvers and the refinement of precise postures. An asymmetric actor critic architecture accelerates training by leveraging privileged simulation information, while noise-injected observations improve robustness under uncertainty. We further show that synergistic wheel–leg coordination reduces joint torque consumption by $15.8\%$ and $26.2\%$ and improves stabilization through energy transfer mechanisms. Extensive evaluations on two distinct quadruped platforms achieve recovery success rates up to $99.1\%$ and $97.8\%$ without platform-specific tuning. Supplementary material is available at https://anonymous.4open.science/w/L2R-WheelLegCoordination-1A73/
@@ I Introduction
Wheeled-legged robots, with their integrated wheel–leg structure, have demonstrated unique mobility advantages in complex environments, establishing them as key platforms for long-duration tasks such as inspection and exploration[1, 2, 3, 4]. However, unexpected falls resulting from dynamic disturbances (e.g. collisions and slippage) often interrupt missions, and existing systems typically rely on manual intervention for recovery or preplanned recovery motions, which, while they can be effective on flat terrains, are not robust against diverse terrain geometries, severely limiting operational efficiency and autonomy. Thus, breakthroughs in robust and autonomous post-fall recovery are critical to achieving fully autonomous missions [5].
@@
In essence, recovery after a fall is a highly complex motion planning and control problem, necessitating precise posture adjustments through multiple discontinuous contact events[6], while the ground specific geometry can impose additional contact uncertainty challenges, which can compromise the successful execution of the fall recovery actions. For the controller, deriving an optimal or near-optimal sequence of actions is highly challenging. Due to the highly nonlinear dynamics of wheeled-legged robots, optimization-based methods often depend on simplified system models and precise state estimation. Consequently, most of the existing work has focused on pure legged recovery[7] mainly demonstrated on flat terrains, leaving the full potential of wheel–leg collaborative recovery largely unexplored. In contrast, recent advances in reinforcement learning have demonstrated remarkable progress in addressing this issue, offering a new technical pathway toward more flexible and robust recovery strategies.
Figure 1: Wheel and joint coordination for joint reset and Center-of-Mass adjustment during the recovery process.
@@
In this paper, we explore a widely used learning-based approach to facilitate rapid post-fall recovery for wheeled-legged robots by leveraging their wheels. While most existing approaches struggle with sparse rewards, unstable convergence, and poor transferability across platforms, our method introduces an asymmetric Proximal Policy Optimization (PPO) framework that integrates dynamic episode exploration and curriculum learning. This design mitigates the sensitivity of recovery and smoothness weight parameters while also avoiding an overly conservative policy tendency. The main contributions of this work are summarized as follows:
@@
  • Adaptive Learning Framework: We propose an Episode-based Dynamic Reward Shaping mechanism combined with curriculum learning to balance exploration and policy convergence. This framework facilitates broad discovery of diverse recovery maneuvers in early training and gradually prioritizes precise posture control as learning progresses. It effectively mitigates the pitfalls of sparse reward formulations and unstable convergence common in prior reinforcement learning (RL)-based recovery methods.
@@
  • Quantified Wheel–Leg Synergy: We systematically validate the critical role of wheel actuation in reducing joint torque effort and improving stabilization during recovery. This synergy between wheel dynamics and leg adjustments is rigorously quantified through cross-platform experiments, offering new insights into how wheels can actively assist in upright recovery processes.
@@
  • Cross-Platform Generalization: We conduct extensive experimental validation on two distinct wheeled-legged quadrupeds (KYON and Unitree Go2-W[8]) demonstrating high adaptability to diverse hardware configurations. The proposed framework achieves recovery success rates above $97\%$ and substantial torque reduction without platform-specific tuning, underscoring its robustness, scalability, and real-world deployability.
@@ II RELATED WORK
In reinforcement learning–based recovery, using only the terminal posture as a reward signal results in sparse and delayed feedback, which hinders exploration and limits policy robustness. To overcome this, we propose an Episode-based Dynamic Reward Shaping mechanism that adapts reward composition throughout training to balance exporation and convergence. This approach promotes broad discovery of recovery maneuvers in the early stages, while gradually and gradually emphasizes precise posture refinement ensuring that the policy effectively converges to the desired recovery behavior.
Figure 2: (a)-(c) correspond to the initial states of different episodes. KYON Legged-wheeled Robot model: The presented model corresponds to a new robot under development.
@@ III-A State initialization and rollout
During the execution of various tasks, wheeled-legged robots may fall or even completely roll over; thus, the recovery controller must accommodate various initial conditions. To simulate different fallen states, we randomly initialize the robot’s base orientation and joint angles, set the joint torques to zero, and let the robot free-fall from a height of 1.1 m for 2 seconds. This duration is chosen to ensure that the robot fully collapses onto the ground, as shown in Fig. 2, which helps avoid the situation where the dynamics become unrealistic[12]. The diverse set of observed states improves policy generalization and adaptability. Each episode has a fixed duration of 5 seconds and is not terminated early.
@@ III-B Privileged information based Actor-Critic
Actor Observation: The policy network observes the robot’s state, including current and historical readings (with a time interval of 0.01 s) of the base’s linear velocity, angular velocity, orientation, joint positions, joint velocities, and wheel speeds. This temporal information enables the robot to better evaluate its contact state[12]. The actor also receives the previous action taken. Except for the most recent action and joint positions, Gaussian noise is added to the observations to account for state-estimation inaccuracies that arise after a fall. Similarly to[5], we apply higher noise levels specifically to joint velocities and the base’s linear and angular velocities.
@@ III-D Dynamic Exploration strategy
1) Dynamic Reward Shaping:In previous research, the post-fall recovery process has often been simplified to a “final pose pursuit” task, typically employing relatively fixed reward or penalty terms based on joint angles, body height, or orientation error. However, in transitioning from a prone position to standing, a wheeled-legged robot may need to undergo large—and sometimes counterintuitive—rolling or swinging movements. Penalizing only the terminal pose too strongly can restrict exploration of diverse state-action pairs, as illustrated in Fig. 2. To balance free exploration in the early stage with precise posture control in the later stage, we propose Episode-based Dynamic Reward Shaping(ED), formally defined as:
$ed=\left(\frac{a\cdot t}{T}\right)^{k},t\in[0,T]$ (1)
where $t$ is the current step in an episode. $T$ is the total number of steps per episode, and $k$ serves as the growth rate, whereas $a$ is the baseline coefficient. For our experiments, we set $a=\frac{T}{2}$, according to the expected recovery time, $k=3$. The immediate reward at each time step is given by:
$r_{t}=ed\cdot R\left(s_{t},a_{t}\right)$ (2)
Here, $R(s_{t},a_{t})$ is the original task-related reward function. When the $t$ is close to 0, the agent can attempt substantial posture adjustments, and any successful state-action pairs are retained and reinforced in the policy network. As $t$ approaches the maximum length of the episode, the emphasis is progressively increased on the refinement of the final target pose. This incremental mechanism promotes action diversity in the early stage of each episode and ensures the stability of the standing posture in the later stage.
@@
where $\beta$ is the progress rate. We multiply the behavior reward by this weight and update it before each environment reset. This approach ensures rapid acquisition of task-related behaviors in the early stages and smoother motion in the later stages, striking a balance between efficiency and feasibility.
@@
body collision $\sum_{b\in B}\|\lambda_{b}[t]\|^{2}$ $5\times 10^{-2}$
action rate $\sum_{t}(\mathbf{a}[t]-\mathbf{a}[t-1])^{2}$ $1\times 10^{-2}$
joint velocity $\sum_{j}\dot{\mathbf{q}_{j}}^{2}$ $2\times 10^{-2}$
torques $\sum_{j}\tau_{j}^{2}$ $2.5\times 10^{-5}$
acceleration $\sum_{j}\ddot{\mathbf{q}_{j}}^{2}$ $2.5\times 10^{-7}$
wheel velocity $\sum_{k}\dot{\mathbf{q}_{k}}^{2}$ $2\times 10^{-2}$
Figure 4: Recovery processes under the same initial posture using two different strategies. (a)–(d) shows the ED strategy successfully recovery. (e)–(h) illustrates the baseline strategy, which adjusts joint positions and base height but fails to optimize base orientation, becoming trapped in a local optimum and not fully recovering. Full videos could be found on the webpage.
@@ III-E Reward function
Owing to the introduction of the dynamic exploration mechanism, during the early training stages, the agent initially focuses on utilizing joint motions to reorient the body at the beginning of each episode, and gradually shifts its attention toward refining joint angles and base height in the later phase of the episode.
@@
$r_{t}=$ $\,ed\cdot(r_{joint\_pos}+r_{base\_height}+r_{base\_ori})$
$+cw\cdot\big(r_{collision}+r_{action\_rate}$
$+r_{joint\_vel}+r_{torque}+r_{joint\_acc})$
@@ IV-A1 Simulation Setup
All experiments are conducted in the Isaac Sim simulator using KYON and Unitree Go2-W[8] robot models. To enhance policy robustness, extensive domain randomization is applied: the robot’s base mass is perturbed by $\pm 5.0kg$ at the start of each episode, link masses vary by $\pm 10\%$ from their nominal values. Moreover, contact dynamics are randomized by uniformly sampling static friction coefficients from $[0.7,1.3]$ for all robot bodies, while actuator stiffness and damping are perturbed by $\pm 1.0$ and $\pm 0.1$ respectively. Additionally, action noise characterized by a zero mean and a standard deviation of $0.02$ is incorporated.
@@
Each episode lasts 5 seconds (at a control rate of 50Hz) and terminates solely upon timeout. The training framework employs an asymmetric PPO algorithm with a learning rate of $0.001$ and a discount factor of $\gamma=0.99$. Furthermore, curriculum learning is integrated, beginning with a difficulty factor of $0.3$ that decays exponentially at a rate of $0.968$ to progressively enhance policy exploration.
IV-A2 Sim-to-Real
We deploy and evaluate the learned policy on a Unitree Go2W[8], performing zero-shot sim-to-real transfer. The trained policy is available on the project website.
IV-B Ablation of the Episode-based Dynamic Reward Shaping Mechanism
@@
To verify the effectiveness of the ED, we compared the recovery performance of the ED-Policy with that of a baseline under an identical reward function and hyperparameter framework. Recovery was defined as achieving, at the final time step of an episode, a base height exceeding $0.42m$, a deviation of the joint angle from the default configuration of no more than $0.5rad$, and a maximum joint velocity below $0.1$ rad/s. Testing on $2048$ randomly generated initial conditions, the ED-Policy achieved a success rate of $99.1\%$, representing a $2.7$ percentage point improvement over baseline $96.4\%$.
@@
Further analysis revealed that the failure cases for the baseline were predominantly associated with scenarios in which the front leg joint angles were minimally randomized, shown in Fig. 4. To illustrate this, we selected a representative initial state from one of the baseline’s failure cases and used it as the starting state for the ED-Policy to compare the resulting action sequences. The experiments indicate that, due to the penalty imposed on joint angle deviations, the baseline suppressed front leg motions and attempted to adjust the base height solely through the hind legs, ultimately becoming trapped in a local optimum.
@@
This phenomenon is directly related to differences in the exploration capacity of the action space. Principal Component Analysis (PCA) was applied to the action sequences from the 2048 experiments, and the resulting visualization Fig. 5 shows that the variance along the principal components for the ED-Policy ($\sigma^{2}_{PC1}=20.91$, $\sigma^{2}_{PC2}=9.72$) is significantly higher than that for the baseline ($\sigma^{2}_{PC1}=17.33$, $\sigma^{2}_{PC2}=5.22$). These findings confirm that the dynamic reward shaping mechanism, by expanding the exploration boundaries and generating less conservative action sequences, enhances recovery robustness in complex scenarios.
@@
Figure 5: PCA-based comparison of single-episode action distributions for ED-policy and baseline-policy across 2048 environments.
@@ IV-C Wheel-Assisted Recovery
To evaluate the dynamic contribution of wheeled actuation during recovery, we compared the performance differences between a wheel–leg collaborative mode and a pure legged mode. In the experimental setup, the pure legged group was configured by modifying the robot description file to change the wheel joints from continuously actuated to fixed constraints, and by removing the wheel speed penalty term to eliminate policy bias; all other hyperparameters and reward functions were kept identical to those of the wheel–leg collaborative group. Testing across 2048 random initial states demonstrated that the wheel–leg collaborative mode significantly enhanced recovery stability and optimized joint torque distribution. As shown in Fig. 6, although both strategies’ base height trajectories converged to the target height ($0.42m$) within $0.7s$, the wheel–leg collaborative group exhibited lower variance during both the recovery phase ($t<0.7s$) and the stabilization phase ($0.7s<t<1.5s$). In contrast, the pure legged group showed larger fluctuations during the stabilization phase, with a variance of $0.102$ at $t=1s$. This indicates that the base height recovery process was considerably more stable, demonstrating that wheel-assisted rolling effectively mitigates the body oscillations observed in the pure legged recovery.